OpenStudy (anonymous):

Find the distance for the following two points A (-10,-15) (-5, -3).

OpenStudy (anonymous):

13 units

OpenStudy (anonymous):

\[\sqrt{(x1-x2)^2+(y1-y2)^2}=\sqrt{(-10+5)^2+(-15+3)^2}=\sqrt{25+144}\]